# Quartet: Native FP4 Training Can Be Optimal for Large Language Models

### v3

@@ 4.2 Ingredient 2: Mixed-Precision Induces Inference-Training Trade-Offs
The above scaling law suggests that, given a set of scaling parameters and a target loss we wish the model to achieve, we can directly solve for the “optimal” forward and backward precisions which allow us to match the loss. However, as pointed out by sardana2025chinchillaoptimalaccountinginferencelanguage, it is often the case in practice that we wish to put a larger weight on inference cost, rather than training cost, which can lead to different results when determining the “optimal” training precisions. Because inference latency depends solely on the forward pass ($\sim 33\%$ of training compute) while the backward pass consumes the remaining $\sim 66\%$, these trade-offs may need to be analyzed separately.
@@
We contrast this with previous work, which often requires lower precision to suffer no accuracy loss (e.g., chmiel2024accurateneuraltraining4bit). This unnecessarily reduces these trade-offs to simple selection of the fastest lossless precision. We argue that scaling-law analysis enables a more fine-grained approach needed to decide upon the “optimal” set of forward and backward precisions.
@@
4.3 Ingredient 3: Minimal Forward-Pass Error and Unbiased Gradient Estimation
The above ingredients should allow us to determine the “best” quantized training method among existing approaches, focusing on the hardware-supported MXFP4 [mxfp] format.
@@ 4.3 Ingredient 3: Minimal Forward-Pass Error and Unbiased Gradient Estimation
As detailed in Section 2, existing QAT (forward-only) approaches can be split into “noise injection” [baskin2021uniq] and “error-minimization” approaches, e.g. [panferov2025queststabletrainingllms]. Focusing on the forward pass, by the above discussion (Ingredients 1 and 2), we seek the approach which maximizes the parameter efficiency factor $\text{eff}_{N}$. For this, we implement four standard schemes for QAT: 1) stochastic rounding (SR) with standard AbsMax per-group normalization [tseng2025trainingllmsmxfp4]; 2) vanilla round-to-nearest (RTN) quantization with AbsMax per-group normalization; 3) learnable scale clipping (LSQ) with RTN quantization [esser2019learned; xi2023int4]; 4) Hadamard normalization followed by RMSE-based clipping (QuEST) [panferov2025queststabletrainingllms]. For fairness, we apply the Hadamard transform to weights and activations for each one of these schemes before quantization. We compare these approaches following Section 4.1: we train models using each technique, apply scaling law fitting, and register their resulting $\text{eff}_{N}$ factors. For additional information, we also show representations’ mean-squared error (MSE) for fitting random Gaussian data. The results are provided in the first rows/columns of Table 2.
Table 2: Illustration of error-bias trade-off between different quantized forward and backward pass approaches. For the forward (given by the $\text{eff}_{N}$ metric) the best performing method is QuEST, correlating with superior MSE over Gaussian input data. By contrast, for the backward pass (the data efficiency $\text{eff*}_{D}$ computed at 800 Tokens/Parameter), the best performing method is stochastic rounding, correlated with perfect magnitude alignment. This justifies our choice of method, which combines block-wise QuEST on the forward, with Stochastic Rounding on the backward pass.
@@
Stochastic Rounding AbsMax 0.44 $2.77\times 10^{-2}$ 0.85 0
Round-to-nearest AbsMax 0.61 $1.37\times 10^{-2}$ 0.83 $9.3\times 10^{-3}$
QuEST (Hadamard + RMSE) 0.65 $1.32\times 10^{-2}$ 0.18 $1.3\times 10^{-2}$
The results in Table 2 show that QuEST has the best parameter efficiency $\text{eff}_{N}$ among all existing methods. Moreover, $\text{eff}_{N}$ appears to correlate heavily with MSE, as suggested by panferov2025queststabletrainingllms; panferov2025unifiedscalinglawscompressed. Additionally, the results align with the analysis of chmiel2024accurateneuraltraining4bit that determined deterministic RTN to always be preferable to stochastic rounding for the forward pass.
@@
The above findings do not transfer to backward pass quantization, as optimization theory shows that unbiased gradient estimation is critical for convergence, e.g. [alistarh2017qsgd]. This leads to a trade-off between the error minimization we can obtain on the forward pass, and the bias induced over the backward pass for a given method. We study this trade-off via a novel analysis of gradient alignment between different quantization methods.
To study gradient bias, we follow the analysis of [vargaftik2021driveonebitdistributedmean; vargaftik2022edencommunicationefficientrobustdistributed], who studied RTN quantization with randomized rotations, approximated by the randomized Hadamard transform, which we denote by $\widehat{H}$. They show that, while RHT makes quantization unbiased in direction, it adds a bias in magnitude. To address this, they proposed an approach that makes RTN projections of post-RHT vectors unbiased, denoted by $Q$, via the following input ($X$) and randomness ($\xi$) specific group-wise rescaling factor $S$:
@@
Unfortunately, their re-scaling is incompatible with coarse group-wise scaling of the MXFP4 format, so we cannot use it in practice. However, we can still use their approach to gauge the degree of misalignment for different quantizers by simply studying their corresponding expected value of $1-\mathbb{E}\left[1/S\right]$, which we call the projection magnitude misalignment. This factor is presented in Table 2, along with the MSE across different schemes. Focusing on stochastic rounding (SR) vs round-to-nearest (RTN) with AbsMax, one can see that SR trades high error for perfect alignment.
@@
Interestingly, we observe that cosine similarity (and MSE by extension) has a high impact on initial convergence and shorter training runs, while projection magnitude misalignment has greater impact on longer runs. Concretely, while RTN backward quantization may be preferable for shorter training, stochastic rounding (SR) performs consistently better for models more saturated with data. In this setup, the inflection point is around the $D/N=400$ data-to-parameter ratio.
@@
Our analysis outlines a new trade-off between parameter efficiency on the forward (equated with quantization MSE), and data-efficiency on the backward (which we equate with the new misalignment metric). In the following, we will adopt a “best of both worlds” approach, aiming to perform a forward pass that minimizes MSE (based on QuEST [panferov2025queststabletrainingllms]) together with a backward pass that is unbiased (based on Stochastic Rounding [tseng2025trainingllmsmxfp4]). The novel challenge, which we address next, will be an extremely efficient GPU-aware implementation of such an approach.
@@ 4.4 Ingredient 4: Fast GPU Support for Accurate Quantized Training
4:  $G_{q}\leftarrow\mathrm{SR}(\frac{3}{4}G_{{h}})$;  $W_{q}^{\top}\leftarrow\mathrm{SR}(\frac{3}{4}W_{{h}}^{\top})$
@@
8:  ${G}^{\top}_{q}\leftarrow\mathrm{SR}(\frac{3}{4}G^{\top}_{h})$;  $X^{\top}_{q}\leftarrow\mathrm{SR}(\frac{3}{4}X^{\top}_{h})$
@@
Quartet Overview.
We integrate our prior discussion into Algorithm 1, which aims to perform accurate training while executing all three matrix multiplications of a linear layer in low precision. The forward pass applies a fixed Hadamard transform $\mathrm{H}_{g}$ (of block size $g$ equal to the quantization group size) and QuEST projection to low precision and multiplies the resulting tensors with an MXFP44 kernel. The backward pass decorrelates the multiplied tensors with an identical block‑wise random Hadamard transform $\mathrm{\widehat{H}}_{g}$, applies unbiased stochastic rounding (SR) to MXFP4, performs the two gradient GEMMs in MXFP4, rescales to compensate for SR range matching, applies QuEST masks ($M_{x}$$M_{w}$) and inverts the Hadamard transform $\mathrm{H}_{g}$.
Costs and format specialization.
GPU kernel support.
While the above blueprint appears simple, implementing it efficiently on Blackwell GPUs—in order to leverage fast MXFP4 support—is extremely challenging. For illustration, a direct implementation of the above pattern would be slower than FP16 unquantized training, let alone optimized FP8. Our fast implementation builds on CUTLASS 3.9 [Thakkar_CUTLASS_2023], which provides templates for the new Blackwell architecture. Computation happens in two stages: Stage 1 fuses the Hadamard transform, quantization, scale calculation, and QuEST clipping mask generation (only on forward) into a single kernel; Stage 2 performs GEMM using a dedicated kernel.
Stage 2: Dedicated GEMM kernel. Blackwell introduces the tcgen05.mma instructions, which natively support matrix multiplication with scale factors in the form $D=C+(A\times\mathrm{SFA})\cdot(B\times\mathrm{SFB})$. These scale factors are applied along the inner ($K$) dimension of the GEMM. For MXFP types, every 32 elements along the $K$-dimension of matrices $A$ and $B$ share a corresponding scale factor. This implies that an $M\times K$ matrix $A$ is associated with a scale matrix $\mathrm{SFA}$ of size $M\times\left\lceil K/32\right\rceil$. Our dedicated kernel is based on CUTLASS block-scaled GEMM for narrow precision. As part of this implementation, we also included the necessary functions to reorganize the scale factors generated in the Stage 1, aligning them with the layout required by this architecture [Blackwell].
5 Experiments
We now provide additional experimental support for the validity of Quartet, focusing on accuracy comparisons with existing INT4/FP4 training methods, and examining kernel speedups.
Experimental setup and scaling law fit.
As described in Section 3, we pre‑train Llama‑style models on C4 and report validation loss after a fixed token budget. All baselines reuse the optimizer, schedule, and hyper‑parameters, as described in Appendix A.1. Following Section 4.1, we compare accuracy across methods by fitting the full scaling law in Eqn. 1 across methods, as follows: we fit parameters $A,\alpha,B,\beta,E$ and $\gamma$ on a grid of baseline precision runs (FP8 forward, FP8 backward) shown on Figure 1(a). Then we fit the parameter and data efficiencies $\text{eff}_{N}$ and $\text{eff}_{D}$ separately for every forward and backward quantization scheme we evaluate. The law is fitted identically to prior work in this area [hoffmann2022trainingcomputeoptimallargelanguage; kumar2024scalinglawsprecision; busbridge2025distillationscalinglaws]. For a more detailed description we refer to Appendix A.2.
Accuracy comparisons.
We compare accuracy (validation loss) as well as the efficiency factors against four recent, fully–quantized training pipelines that operate in 4‑bit precision for both forward and backward passes: 1) LUQ [chmiel2024accurateneuraltraining4bit] applies to both INT4 and FP4, using unbiased quantization that pairs 4-bit weights/activations with stochastic underflow, and logarithmic stochastic rounding; 2) HALO [ashkboos2025halohadamardassistedlowerprecisionoptimization], which uses Hadamard rotations to mitigate outliers, evaluated in FP4 at their most accurate HALO‑2 setting; 3) Jetfire [xi2024jetfire] performs quantization in blocks of $32\times 32$, originally introduced for INT8, and adapted to FP4 for our setup; 4) LSS [xi2023int4] for INT4 training, that combines a Hadamard‑based forward pass with “leverage–score” sampled INT4 gradients.
[…]
LUQ–INT4 3.73 3.68 3.66 3.43 3.40 0.50 0.15
LUQ–FP4 4.81 4.91 4.88 4.84 4.80 0.01 0.09
Jetfire–FP4 7.03 6.94 6.76 6.62 6.58 0.01 0.07
HALO–FP4 6.65 7.04 6.55 6.50 5.38 Unstable
LSS–INT4 NaN 3.40 NaN NaN NaN Unstable
Quartet 3.50 3.38 3.30 3.24 3.21 0.64 0.94
Accuracy discussion.
[…]
Additionally, we trained 100M, 200M, 430M, 800M and 1.6B parameters Llama models with Quartet and FP8, with $D/N=100$. We evaluated them on a set of few-shot benchmarks, including HellaSwag [zellers2019hellaswagmachinereallyfinish], WinoGrande [sakaguchi2019winograndeadversarialwinogradschema] and ARC-easy [clark2018thinksolvedquestionanswering]. Figure 4 demonstrate that those evaluations are consistent with C4 validation loss for larger models.
Speedup results.
6 Discussion and Limitations
We provided a set of guidelines to modeling, comparing and designing fully-quantized training schemes for large language models. Moreover, we followed those guidelines to arrive at Quartet: a new SOTA full MXFP4 training algorithm. One current limiting factor is that Quartet was designed with a specific (standard) data-type and compute architecture in mind. Certain aspects of our method rely on specialized operations, like stochastic rounding, which have hardware support for MXFP4, but may be lacking for other formats. In future work, we plan to look into generalizing our approach to alternative formats, as well as larger-scale distributed model execution.
Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12, i.e., the Bilateral AI Cluster of Excellence, and through generous gifts by NVIDIA and Google.
Appendix A Technical Appendices and Supplementary Material
A.1 Training Hyper-parameters

### v4

@@ 4.2 Ingredient 2: Mixed-Precision Induces Inference-Training Trade-Offs
The above scaling law suggests that, given a set of scaling parameters and a target loss we wish the model to achieve, we can directly solve for the “optimal” forward and backward precisions which allow us to match the loss. However, as pointed out by Sardana et al. [41], it is often the case in practice that we wish to put a larger weight on inference cost, rather than training cost, which can lead to different results when determining the “optimal” training precisions. Because inference latency depends solely on the forward pass ($\sim 33\%$ of training compute) while the backward pass consumes the remaining $\sim 66\%$, these trade-offs may need to be analyzed separately.
@@
We contrast this with previous work, which often requires lower precision to suffer no accuracy loss (e.g., Chmiel et al. [12]). This unnecessarily reduces these trade-offs to simple selection of the fastest lossless precision. We argue that scaling-law analysis enables a more fine-grained approach needed to decide upon the “optimal” set of forward and backward precisions.
@@
4.3 Ingredient 3: Minimal Forward-Pass Error and Error-Bias Trade-off
The above ingredients should allow us to determine the “best” quantized training method among existing approaches, focusing on the hardware-supported MXFP4 [36] format.
@@ 4.3 Ingredient 3: Minimal Forward-Pass Error and Error-Bias Trade-off
As detailed in Section 2, existing QAT (forward-only) approaches can be split into “noise injection” [5] and “error-minimization” strategies, e.g. [37]. Focusing on the forward pass, by the above discussion (Ingredients 1 and 2), we seek the approach which maximizes the parameter efficiency factor $\text{eff}_{N}$. For this, we implement four standard schemes for QAT: 1) stochastic rounding (SR) with standard AbsMax per-group normalization [47]; 2) vanilla round-to-nearest (RTN) quantization with AbsMax per-group normalization; 3) learnable scale clipping (LSQ) with RTN quantization [20; 58]; 4) Hadamard normalization followed by RMSE-based clipping (QuEST) [37]. For fairness, we apply the Hadamard transform to weights and activations for each one of these schemes before quantization. We compare these approaches following Section 4.1: we train models using each technique, apply scaling law fitting, and register their resulting $\text{eff}_{N}$ factors. For additional information, we also show representations’ mean-squared error (MSE) for fitting random Gaussian data. The results are provided in the first rows/columns of Table 2.
Table 2: Illustration of error-bias trade-off between different quantized forward and backward pass approaches. For the forward (given by the $\text{eff}_{N}$ metric) the best performing method is QuEST, correlating with superior MSE over Gaussian input data. By contrast, for the backward pass (the data efficiency $\text{eff}_{D}$), the best performing method is RTN, that balances quadratic error with magnitude alignment. This justifies our choice of method, which combines block-wise QuEST on the forward, with RTN on the backward pass.
@@
Stochastic Rounding AbsMax 0.42 $2.77\times 10^{-2}$ 0.88 0
Round-to-nearest AbsMax 0.59 $1.37\times 10^{-2}$ 0.93 $9.3\times 10^{-3}$
QuEST (Hadamard + RMSE) 0.64 $1.32\times 10^{-2}$ 0.83 $1.3\times 10^{-2}$
The results in Table 2 show that QuEST has the best parameter efficiency $\text{eff}_{N}$ among all existing methods. Moreover, $\text{eff}_{N}$ appears to correlate heavily with MSE, as suggested by Panferov et al. [37, 38]. Additionally, the results align with the analysis of Chmiel et al. [12] that determined deterministic RTN to always be preferable to stochastic rounding for the forward pass.
@@
The above findings do not transfer to backward pass quantization, as optimization theory shows that unbiased gradient estimation is critical for convergence, e.g. [1]. This leads to a trade-off between the error minimization we can obtain on the forward pass, and the bias induced over the backward pass for a given method. We study this trade-off via a novel analysis of gradient alignment between different quantization methods.
To study gradient bias, we follow the analysis of [48; 49], who studied RTN quantization with randomized rotations, approximated by the randomized Hadamard transform, which we denote by $\widehat{H}$. They show that, while RHT makes quantization unbiased in direction, it adds a bias in magnitude. To address this, they proposed an approach that makes RTN projections of post-RHT vectors unbiased, denoted by $Q$, via the following input ($X$) and randomness ($\xi$) specific group-wise rescaling factor $S$:
@@
Unfortunately, their re-scaling is incompatible with coarse group-wise scaling of the MXFP4 format, so we cannot use it in practice. However, we can still use their approach to gauge the degree of misalignment for different quantizers by simply studying their corresponding expected value of $1-\mathbb{E}\left[1/S\right]$, which we call the projection magnitude misalignment. This factor is presented in Table 2, along with the MSE across different schemes. Focusing on stochastic rounding (SR) vs round-to-nearest (RTN) with AbsMax, one can see that SR trades higher error for perfect alignment.
@@
We observe that RTN backward quantization consistently outperforms the alternatives, although the gap with stochastic rounding (SR) decreases as the training duration grows  [^1] In the first draft of the paper, an implementation bug caused large performance degradation for RTN backward quantization for longer training runs. This degradation led to our arriving at a different conclusion (SR superiority), which we now reject..
@@
Our analysis outlines a new trade-off between parameter efficiency on the forward (equated with quantization MSE), and data-efficiency on the backward (which characterize with with a trade-off between error and misalignment). In the following, we will adopt a “best of both worlds” approach, aiming to perform a forward pass that minimizes MSE (based on QuEST [37]) together with an RTN backward pass that is balances quadratic error with and magnitude alignment. The novel challenge, which we address next, will be an extremely efficient GPU-aware implementation of such an approach.
@@ 4.4 Ingredient 4: Fast GPU Support for Accurate Quantized Training
Figure 3: Quartet computation flow for a fully-MXFP4 linear layer training.
Quartet Overview.
We integrate our prior discussion into Algorithm 1 (also shown in Figure 3), which aims to perform accurate training while executing all three matrix multiplications of a linear layer in low precision. The forward pass applies a fixed Hadamard transform $\mathrm{H}_{g}$ (of block size $g$ equal to the quantization group size) and QuEST projection to low precision and multiplies the resulting tensors with an MXFP44 kernel. The backward pass decorrelates the multiplied tensors with an identical block‑wise random Hadamard transform $\mathrm{\widehat{H}}_{g}$, applies round-to-nearest (RTN) to MXFP4, performs the two gradient GEMMs in MXFP4, rescales to compensate for RTN range matching [47], applies QuEST masks ($M_{x}$$M_{w}$) and inverts the Hadamard transform $\mathrm{H}_{g}$.
Costs and format specialization.
GPU kernel support.
While the above blueprint appears simple, implementing it efficiently on Blackwell GPUs—in order to leverage fast MXFP4 support—is extremely challenging. For illustration, a direct implementation of the above pattern would be slower than FP16 unquantized training, let alone optimized FP8. Our fast implementation builds on CUTLASS 3.9 [45], which provides templates for the new Blackwell architecture. Computation happens in two stages: Stage 1 fuses the Hadamard transform, quantization, scale calculation, and QuEST clipping mask generation (only on forward) into a single kernel; Stage 2 performs GEMM using a dedicated kernel.
Stage 2: Dedicated GEMM kernel. Blackwell introduces the tcgen05.mma instructions, which natively support matrix multiplication with scale factors in the form $D=C+(A\times\mathrm{SFA})\cdot(B\times\mathrm{SFB})$. These scale factors are applied along the inner ($K$) dimension of the GEMM. For MXFP types, every 32 elements along the $K$-dimension of matrices $A$ and $B$ share a corresponding scale factor. This implies that an $M\times K$ matrix $A$ is associated with a scale matrix $\mathrm{SFA}$ of size $M\times\left\lceil K/32\right\rceil$. Our dedicated kernel is based on CUTLASS block-scaled GEMM for narrow precision. As part of this implementation, we also included the necessary functions to reorganize the scale factors generated in the Stage 1, aligning them with the layout required by this architecture [35].
5 Experiments
We now provide additional experimental support for the validity of Quartet, shown in Figure 3, focusing on accuracy comparisons with existing INT4/FP4 training methods, and examining kernel speedups.
Experimental setup and scaling law fit.
As described in Section 3, we pre‑train Llama‑style models on C4 and report validation loss after a fixed token budget. All baselines reuse the optimizer, schedule, and hyper‑parameters, as described in Appendix A.2. Following Section 4.1, we compare accuracy across methods by fitting the full scaling law in Eqn. 1 across methods, as follows: we fit parameters $A,\alpha,B,\beta,E$ and $\gamma$ on a grid of baseline precision runs (FP8 forward, FP8 backward) shown on Figure 1(a). Then we fit the parameter and data efficiencies $\text{eff}_{N}$ and $\text{eff}_{D}$ separately for every forward and backward quantization scheme we evaluate. The law is fitted identically to prior work in this area [27; 30; 8]. For a more detailed description we refer to Appendix A.3.
Accuracy comparisons.
We compare accuracy (validation loss) as well as the efficiency factors against four recent, fully–quantized training pipelines that operate in 4‑bit precision for both forward and backward passes: 1) LUQ [12] applies to both INT4 and FP4, using unbiased quantization that pairs 4-bit weights/activations with stochastic underflow, and logarithmic stochastic rounding; 2) HALO [3], which uses Hadamard rotations to mitigate outliers, evaluated in FP4 at their most accurate HALO‑2 setting; 3) Jetfire [57] performs quantization in blocks of $32\times 32$, originally introduced for INT8, and adapted to FP4 for our setup; 4) LSS [58] for INT4 training, that combines a Hadamard‑based forward pass with “leverage–score” sampled INT4 gradients.
[…]
LUQ–INT4 3.73 3.68 3.66 3.43 3.40 0.49 0.15
LUQ–FP4 4.81 4.91 4.88 4.84 4.80 0.01 0.07
Jetfire–FP4 7.03 6.94 6.76 6.62 6.58 Unstable
HALO–FP4 6.65 7.04 6.55 6.50 5.38 Unstable
LSS–INT4 NaN 3.40 NaN NaN NaN Unstable
Quartet 3.49 3.38 3.29 3.24 3.20 0.65 0.95
Accuracy discussion.
[…]
Additionally, we trained 100M, 200M, 430M, 800M and 1.6B parameters Llama models with Quartet and FP8, with $D/N=100$. We evaluated them on a set of few-shot benchmarks, including HellaSwag [60], WinoGrande [40] and ARC-easy [15]. Figure 5 demonstrate that those evaluations are consistent with C4 validation loss for larger models.
Speedup results.
6 Discussion and Limitations
We provided a set of guidelines to modeling, comparing and designing fully-quantized training schemes for large language models. Moreover, we followed those guidelines to arrive at Quartet: a new SOTA full MXFP4 training algorithm. One current limiting factor is that Quartet was designed with a specific (standard) data-type and compute architecture in mind. In future work, we plan to look into generalizing our approach to alternative formats, as well as larger-scale distributed model execution.
Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12, i.e., the Bilateral AI Cluster of Excellence, and through generous gifts by NVIDIA and Google. This work was supported under project ID 40 as part of the Swiss AI Initiative, through a grant from the ETH Domain and computational resources provided by the Swiss National Supercomputing Centre (CSCS) under the Alps infrastructure. We would like to thank our contacts at Datacrunch/Verda (Paul Chang and Antonio Dominguez) for hardware support that was essential to this project.
Appendix A Technical Appendices and Supplementary Material
A.1 Algorithm
@@
4:  $G_{q}\leftarrow\mathrm{RTN}(\frac{3}{4}G_{{h}})$;  $W_{q}^{\top}\leftarrow\mathrm{RTN}(\frac{3}{4}W_{{h}}^{\top})$
@@
8:  ${G}^{\top}_{q}\leftarrow\mathrm{RTN}(\frac{3}{4}G^{\top}_{h})$;  $X^{\top}_{q}\leftarrow\mathrm{RTN}(\frac{3}{4}X^{\top}_{h})$
@@
A.2 Training Hyper-parameters